If the cos of an angle is .75, what is the csc?

Answer by stanbon(75887)

If the cos of an angle is .75, what is the csc?
--------
cos = x/r ; sin = y/r
Your Problem:
cos = x/r=3/4
If x=3 and y=4, r= sqrt(3^2+4^2)=5
-------
So, sin = y/r = 4/5
==========
Cheers,
Stan H.